# DiffEyeSyn: Diffusion-based User-specific Eye Movement Synthesis

### v1

@@ Abstract.
High-frequency components in eye gaze data contain user-specific information promising for various applications, but existing gaze modelling methods focus on low frequencies of typically not more than 30 Hz. We present DiffEyeSyn – the first computational method to synthesise high-frequency gaze data, including eye movement characteristics specific to individual users. The key idea is to consider the high-frequency, user-specific information as a special type of noise in eye movement data. This perspective reshapes eye movement synthesis into the task of injecting this user-specific noise into any given eye movement sequence. We formulate this injection task as a conditional diffusion process in which the synthesis is conditioned on user-specific embeddings extracted from the gaze data using pre-trained models for user authentication. We propose user identity guidance  – a novel loss function that allows our model to preserve user identity while generating human-like eye movements in the spatial domain. Experiment results on two public high-frequency eye movement biometric datasets show that our synthetic eye movements are indistinguishable from real human eye movements. Furthermore, we demonstrate that DiffEyeSyn can be used to synthesise eye gaze data at scale and for different downstream tasks, such as gaze data imputation and gaze data super-resolution. As such, our work lays the methodological foundations for personalised eye movement synthesis that has significant application potential, such as for character animation, eye movement biometrics, or gaze-based activity and context recognition.
ccs: Computing methodologies Artificial intelligenceccs: Human-centered computingccs: Computing methodologies Machine learning
@@ 1. Introduction
While low-resolution eye gaze recordings (typically under 30 Hz) can fulfil the requirements of some applications like gaze-based activity recognition (Bulling et al., 2013; Hu et al., 2023), high-resolution eye gaze signals are key to numerous applications such as fast eye movement detection (Leube and Rifai, 2017; Funke et al., 2016; Dalmaijer, 2014) that detects subtle eye movements like microsaccades from high-resolution gaze recordings as well as gaze-based interaction (Duchowski, 2018; Angelopoulos et al., 2021) that reduces the system delay using high-frequency eye gaze data. More importantly, recent studies in eye tracking research have revealed that the high-frequency components in eye gaze signals contain rich user-specific information (Holland and Komogortsev, 2012; Jiao et al., 2023) that is significant for modelling eye movement behaviours of individual users and has great relevance with various applications including gaze data imputation and gaze data super-resolution. Despite the significance of high-frequency eye gaze data, existing works on modelling eye movement behaviours have only focused on low-frequency eye gaze signals.
In this work, we present DiffEyeSyn – the first computational method to synthesise high-frequency gaze data, including eye movement characteristics specific to individual users. The key idea of our approach is to inject high-frequency, user-specific information into any given eye movement sequence. We formulate this injection task as a conditional diffusion process. The synthesis is conditioned on user-specific embeddings extracted from the gaze data using pre-trained user authentication models. We propose user identity guidance - a novel loss function that allows our model to preserve user identity while generating human-like eye movements in the spatial domain. Experiments on two public eye movement biometric datasets show that our synthetic eye movements are indistinguishable from real human eye movements. Furthermore, we demonstrate that DiffEyeSyn can be used to synthesise eye tracking data at scale and for different downstream tasks, such as gaze data imputation and gaze data super-resolution.
@@
    We propose DiffEyeSyn – a novel diffusion-based method that injects high-frequency, user-specific information into any given eye movement sequence to synthesise high-frequency, user-specific eye movement signals.
@@
    We report extensive experiments on two public eye movement biometric datasets and demonstrate that our synthetic eye movements are indistinguishable from real human eye movement signals.
@@
    We demonstrate our method’s effectiveness in the sample downstream tasks of gaze data imputation and gaze data super-resolution.
@@ 2.1. Eye Movement Synthesis
Recent works have started to incorporate deep learning methods to generate eye movement data. For example, Fuhl et al. (Fuhl et al., 2021) leveraged a variational autoencoder (VAE) to simulate eye-tracking data without specific stimuli. Assens et al. (Assens et al., 2018) used a generative adversarial network (GAN) to simulate fixations and missing saccadic movements. Prasse et al. (Prasse et al., 2023) proposed SP-EyeGAN to synthesise eye movements during reading. SUPREYES (Jiao et al., 2023) aimed at gaze super-resolution, i.e., synthesised high-frequency eye movement data upon existing low-frequency data. Jiao et al. developed a diffusion-based method, DiffGaze, to synthesise eye movements at 30 Hz given the $360^{\circ}$ images in virtual reality (VR) (Jiao et al., 2024).
However, existing methods have only generated ubiquitous scanpaths or low-frequency eye movements, which neglect the user-specific information manifested in high-frequency data. In stark contrast, our DiffEyeSyn is the first approach to synthesise user-specific high-frequency eye movements. In addition, our method can be used in various applications, such as gaze data imputation, gaze data super-resolution, and eye movement generation.
@@ 2.2. Eye Movement Biometrics
These prior works have observed that high-frequency (1000 Hz and higher) components in gaze data contain user-specific information, promising for various applications, such as user authentication (Lohr and Komogortsev, 2022). Our work, for the first time, leverages the pretrained user authentication model for user-specific eye movement synthesis and its corresponding evaluation.
@@ 3. Background: Denoising Diffusion Probabilistic Models
Denoising diffusion probabilistic models (DDPMs) (Ho et al., 2020) is a type of generative model that has demonstrated state-of-the-art performance on various data synthesis tasks, such as image (Dhariwal and Nichol, 2021; Rombach et al., 2021) and video generation (Brooks et al., 2024), audio synthesis (Kong et al., 2020), as well as low-frequency gaze sequence generation (Jiao et al., 2024). DDPMs consist of two computational processes: A forward process and a reverse process. The forward process aims to shift the input data distribution to a normal distribution by adding $T$-step noises. The diffusion model is trained in the forward process to predict how much noise is added to the original data distribution given a noisy distribution and a timestep. The reverse process is where the inference happens. Starting with a normal distribution, the model trained in the forward process gradually predicts the noise for $T$-step to denoise the normal distribution back to the original data distribution.
@@
where $x_{0}$ is the original input data, $x_{t}$ is the noisy data at timestep $t$, and $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ is a hyper-parameter that can be calculated given a predefined noise schedule. The noise schedule indicates the amount of noise should be added at each timestep. Given noisy data $x_{t}$, a deep learning model $f_{\theta}$ is trained to predict the added noise at timestep $t$:
@@
During the reverse process, the generated data is initialised from the normal distribution $x_{t}$, and realistic data is obtained after $T$ denoising steps. The reverse process can be formalised as:
@@ 4.1. Problem Definition
Research on eye movement biometrics (EMB) has indicated the presence of user-specific information within the high-frequency components of eye tracking data (Jiao et al., 2023; Lohr and Komogortsev, 2022; Makowski et al., 2021). It has been shown that computational models for user identification and authentication, when trained with high-frequency eye movement data, achieve the best performance (Lohr and Komogortsev, 2022). Conversely, models that only had access to low-frequency gaze data (30 Hz or lower) have been deemed unsuitable for practical applications (Lohr and Komogortsev, 2022; Jiao et al., 2023). Informed by these findings, we consider user-specific information as a special type of noise in eye movement data. Consequently, we define the task of user-specific eye movement synthesis as the process of injecting user-specific noise into any given eye movement sequence.
@@
In the reverse process, given a target-user embedding and a reference sequence of eye movements, the trained DiffEyeSyn can inject the user-specific information of the target user into the reference eye movements (details in Section 4.5).
@@ 4.2. Conditional Diffusion Model for Eye Movement Synthesis
Furthermore, the reverse process in Equation 4 is extended to:
(8) $p_{\theta}\left(x_{t-1}\mid x_{t}\right)=\mathcal{N}\left(x_{t-1};\mu_{\theta}(x_{t},t\mid(x^{co}_{0},E(x_{0}))),\sigma_{\theta}(x_{t},t\mid(x^{co}_{0},E(x_{0})))I\right)$
@@ 4.3. User Identity Guidance
Training the model with $\mathcal{L}_{noise}$ allows the model to learn the distribution of human eye movements. However, there is no explicit user identity control in the generation process. Inspired by the classifier-guided diffusion model in image generation (Dhariwal and Nichol, 2021), we propose a novel user identity guidance that leverages the pretrained user authenticator $E$ to allow DiffEyeSyn to capture the user-specific information of human eye movements in the forward steps. At timestep $t$ in the forward process, we obtain $x_{t}$ by adding a sampled noise $\epsilon_{t}$ on the target $x_{0}$. We then predict the added noise $\hat{\epsilon_{t}}$ using Equation 6. With the predicted noise $\hat{\epsilon_{t}}$, we can directly estimate the target $\hat{x_{0}}$ by denoising $x_{t}$ using:
@@ 4.4. Data Preprocessing
1:Input: 1,000 Hz gaze data $q$, pretrained user authenticator $E$, total diffusion step $T$
@@
17:     $\nabla_{\theta}({\left\|{\epsilon_{t}-\hat{\epsilon_{t}}}\right\|}+\lambda(1-CosineSimilarity(E(x_{0}),E(\hat{x_{0}})))$
@@ 4.5. Inference
A summary of DiffEyeSyn inference is shown in Algorithm 2. At the inference time, DiffEyeSyn injects the user-specific information into any given 1,000 Hz eye movements. It is worth noting that DiffEyeSyn not only works for user-specific eye movement synthesis but also works for user-specific gaze data imputation and gaze data super-resolution directly. These two tasks aim to fill in missing values in existing eye movement sequences. To directly apply DiffEyeSyn at the inference time to these two tasks. Consider the example of enhancing the resolution of gaze data from 50 Hz to 1,000 Hz. The user may utilise any interpolation technique to up-sample 50 Hz data to 1,000 Hz, which, alongside a sequence of the target user’s eye movements, serves as input to DiffEyeSyn.
@@
1:Input: Any sequence preprocessed reference 1,000 Hz eye movement velocity $x^{co}$, pretrained user authenticator $E$, a sequence preprocessed 1,000 Hz eye movement velocity from the target user $x_{target-user}$, total number of denoising steps $T$
@@
5:      Compute $\mu_{\theta}(x_{t},t\mid(x^{co},E(x_{target-user})))$ and $\sigma_{\theta}(x_{t},t\mid(x^{co},E(x_{target-user})))$ with $\hat{\epsilon_{t}}$
6:      Sample $x_{t-1}\sim p_{\theta}\left(x_{t-1}\mid x_{t}\right)$ using Equation 8
7:end for
8:return $x_{0}$
@@ 4.6. Architecture
(13) $t_{\text{encoding }}=\left[\sin\left(10^{\frac{0\times 4}{63}}t\right),\cdots,\sin\left(10^{\frac{63\times 4}{63}}t\right),\cos\left(10^{\frac{0\times 4}{63}}t\right),\cdots,\cos\left(10^{\frac{63\times 4}{63}}t\right)\right]$
@@ 5.1. Experimental Setup
GazeBase is a large-scale dataset of eye movements collected using an Eyelink 1000 eye tracker that records at 1,000 Hz. It has 12,334 eye tracking records from 322 participants, covering different eye movements across nine rounds of recordings over 37 months. Each round has two sessions, and each session has seven visual tasks: fixations (FXS), horizontal (HSS) and random (RAN) saccades, reading (TEX), watching two videos (VD1 and VD2), and playing gaze-driven games (BLG). Further details on this dataset can be found in GazeBase paper (Griffith et al., 2021). To avoid the information leak, we adopted the dataset split described in the EKYT paper, as the EKYT model was trained on this dataset. In particular, the test set comprises recordings from 59 subjects presenting in round 6, while the training set comprises recordings from the remaining 263 participants. The recordings of the BLG task were excluded from the training set. Refer to the EKYT paper by (Lohr and Komogortsev, 2022) for more details. There is no overlap of users between the training and test sets. These ensure zero information leakage from the pretrained model and we are able to test DiffEyeSyn performance on synthesising user-specific eye movements for unseen users during training.
@@
User Identity Removal. Previous studies in eye movement biometrics (Lohr and Komogortsev, 2022; Lohr et al., 2022) and gaze data super-resolution (Jiao et al., 2023) suggest that low-frequency gaze data at 20-30 Hz performs poorly in gaze-based user identification and authentication. In addition, Jiao et al. (Jiao et al., 2023) demonstrate that upsampling the low-frequency gaze data to a high frequency using traditional interpolation methods leads to a performance drop in user identification compared with the original low-frequency data. Based on these findings, we removed user-specific information in the ground truth gaze data by downsampling and interpolations. Specifically, we first downsampled the 1,000 Hz gaze data to 20 Hz and upsampled it back to 1,000 Hz using the previous interpolation. To verify the effectiveness of our identity removal, we calculated the cosine similarity between the EKYT embeddings extracted from the ground truth gaze data and the gaze data after the identity removal on both datasets. The resulting mean cosine similarity values on both datasets, as shown in Table 1, are consistently below $0.1$, confirming the effectiveness of our approach in removing user-specific information from the gaze data. Additionally, compared with adding random noise to the eye movements, removing the user-specific information by downsampling and interpolations keeps the low-frequency basis of the eye movements, which simplifies the task for DiffEyeSyn as it only needs to inject the high-frequency user-specific information back to the given eye movement sequence, instead of learning to synthesis both realistic low-frequency and high-frequency eye movements at once.

### v2

@@ Abstract.
High-frequency gaze data contains more user-specific information than low-frequency data, promising for various applications. However, existing gaze modelling methods focus on low-frequency data, ignoring user-specific subtle eye movements in high-frequency eye movements. We present DiffEyeSyn – the first computational method to synthesise eye movements specific to individual users. The key idea is to consider the user-specific information as a special type of noise in eye movement data. This perspective reshapes eye movement synthesis into the task of injecting this user-specific noise into any given eye movement sequence. We formulate this injection task as a conditional diffusion process in which the synthesis is conditioned on user-specific embeddings extracted from the gaze data using pre-trained models for user authentication. We propose user identity guidance  – a novel loss function that allows our model to preserve user identity while generating human-like eye movements in the spatial domain. Experiments on two public datasets show that our synthetic eye movements preserve user-specific characteristics and are more realistic than baseline approaches. Furthermore, we demonstrate that DiffEyeSyn can synthesise large-scale gaze data and support various downstream tasks, such as gaze-based user identification. As such, our work lays the methodological foundations for personalised eye movement synthesis that has significant application potential, such as for character animation, eye movement biometrics, and gaze data imputation.
submissionid: 1285journalyear: 2025copyright: acmlicensedconference: ; booktitle: ;price: 15.00submissionid: 1285
Figure 1. DiffEyeSyn synthesises user-specific eye movements at 1,000 Hz (Left), demonstrating the potential for applications such as eye movement biometrics and rendering personalised eye animations (Right, rendered using the tool proposed in (Świrski and Dodgson, 2014), with further examples in the supplementary video).
@@ 1. Introduction
While low-resolution eye gaze recordings (typically under 30 Hz) can fulfil the requirements of some applications like gaze-based activity recognition (Bulling et al., 2013; Hu et al., 2023), high-resolution eye gaze signals are key to numerous applications such as fast eye movement detection (Leube and Rifai, 2017; Funke et al., 2016; Dalmaijer, 2014) that detects subtle eye movements like microsaccades from high-resolution gaze recordings as well as gaze-based interaction (Duchowski, 2018; Angelopoulos et al., 2021) that reduces the system delay using high-frequency eye gaze data. More importantly, recent studies in eye tracking research have revealed that the high-frequency eye gaze signals contain rich user-specific information (Holland and Komogortsev, 2012; Jiao et al., 2023). More importantly, recent studies in eye tracking research have revealed that the high-frequency eye gaze signals contain rich user-specific information (Holland and Komogortsev, 2012; Jiao et al., 2023; Lohr and Komogortsev, 2022; Lohr et al., 2022) than low-frequency ones. For example, the performance of using 60-second 30 Hz eye movements falls far short of the performance using 5-second 1,000 Hz monocular eye movements in gaze-based user authentication (Lohr and Komogortsev, 2022). This indicates that user-specific information within high-frequency eye movement recordings is significant for modelling eye movement behaviours of individual users and has great relevance with various applications, including personalised digital animation (see Figure 1 for example), eye movement biometrics, and gaze data imputation. Despite the significance of high-frequency eye gaze data, existing works on modelling eye movement behaviours have only focused on low-frequency eye gaze signals.
In this work, we present DiffEyeSyn – the first computational method to synthesise high-frequency gaze data, including eye movement characteristics specific to individual users. The key idea of our approach is to inject user-specific information into any given eye movement sequence. We formulate this injection task as a conditional diffusion process. The synthesis is conditioned on user-specific embeddings extracted from the gaze data using pre-trained user authentication models. We propose user identity guidance - a novel loss function that allows our model to preserve user-specific characteristics while generating human-like eye movements in the spatial domain. Experiments on two public eye movement biometric datasets show that our synthetic eye movements contain user-specific information and are more realistic than existing baseline approaches. Furthermore, we demonstrate that DiffEyeSyn can be used to synthesise eye tracking data at scale and for different downstream tasks, such as gaze-based user identification.
@@
    We propose DiffEyeSyn – a novel diffusion-based method that injects user-specific information into any given eye movement sequence to synthesise high-frequency, user-specific eye movement signals.
@@
    We report extensive experiments on two public eye movement biometric datasets and demonstrate that our synthetic eye movements contain user-specific information and are more naturalistic than existing eye movement synthesis methods.
@@
    We demonstrate our method’s effectiveness in the sample downstream task of gaze-based user identification.
@@ 2.1. Eye Movement Synthesis
Recent works have started to incorporate deep learning methods to generate eye movement data. For example, Fuhl et al. (Fuhl et al., 2021) leveraged a variational autoencoder (VAE) to simulate eye-tracking data without specific stimuli. Prasse et al. (Prasse et al., 2023, 2024) proposed SP-EyeGAN to synthesise subtle eye movements in-between the known fixation locations. SUPREYES (Jiao et al., 2023) aimed at gaze super-resolution, i.e., synthesised high-frequency eye movement data upon existing low-frequency data. Jiao et al. developed a diffusion-based method, DiffGaze, to synthesise eye movements at 30 Hz given the $360^{\circ}$ images in virtual reality (VR) (Jiao et al., 2024).
However, existing methods have only generated ubiquitous scanpaths or low-frequency eye movements, which neglect the user-specific subtle eye movements manifested in high-frequency data. In stark contrast, our DiffEyeSyn is the first approach to synthesise user-specific high-frequency eye movements. In addition, a key advantage of DiffEyeSyn is that it can inject these subtle movements into any eye movement sequence, including scanpaths. As such, DiffEyeSyn can serve as a post-processing method for existing stimulus-driven scanpath prediction approaches. Moreover, our method can be used in various applications, such as eye movement biometrics and personalised eye movement animations.
@@ 2.2. Eye Movement Biometrics
These prior works have observed that high-frequency (1000 Hz and higher) gaze data contain more user-specific information, promising for various applications, such as user authentication (Lohr and Komogortsev, 2022). Our work, for the first time, leverages the pretrained user authentication model for user-specific eye movement synthesis and its corresponding evaluation.
@@ 3. Background: Denoising Diffusion Probabilistic Models
Denoising diffusion probabilistic models (DDPMs) (Ho et al., 2020) is a type of generative model that has demonstrated state-of-the-art performance on various data synthesis tasks, such as image (Dhariwal and Nichol, 2021; Rombach et al., 2021) and video generation (Brooks et al., 2024), audio synthesis (Kong et al., 2020), as well as low-frequency gaze sequence generation (Jiao et al., 2024). DDPMs consist of two computational processes: A forward process and a reverse process. The forward process aims to shift the input data distribution to a normal distribution by adding $T$-step noises. The reverse process is where the training and inference happen. The diffusion model is trained in the reverse process to predict how much noise is added to the original data distribution given a noisy distribution and a timestep. Then, starting with a normal distribution, the trained model gradually predicts the noise for $T$-step to denoise the normal distribution back to the original data distribution.
@@
where $x_{0}$ is the original input data, $x_{t}$ is the noisy data at timestep $t$, and $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ is a hyper-parameter that can be calculated given a predefined noise schedule. The noise schedule indicates the amount of noise should be added at each timestep.
During the reverse process, given noisy data $x_{t}$, a deep learning model $f_{\theta}$ is trained to predict the added noise at timestep $t$:
@@
Once the model is trained, the generated data is initialised from the normal distribution $x_{t}$, and realistic data is obtained after $T$ denoising steps. The reverse process can be formalised as:
@@ 4.1. Problem Definition
Research on eye movement biometrics (EMB) has indicated that high-frequency eye tracking data contains more user-specific information than low-freqency eye tracking data. (Jiao et al., 2023; Lohr and Komogortsev, 2022; Makowski et al., 2021). It has been shown that computational models for user identification and authentication, when trained with high-frequency eye movement data, achieve the best performance (Lohr and Komogortsev, 2022). Conversely, models that only had access to low-frequency gaze data (30 Hz or lower) have been deemed unsuitable for practical applications (Lohr and Komogortsev, 2022; Jiao et al., 2023). Informed by these findings, and given higher-frequency data captures more subtle eye movements, we consider user-specific information as a special type of noise in eye movement data. Consequently, we define the task of user-specific eye movement synthesis as the process of injecting user-specific noise (subtle eye movements) into any given eye movement sequence.
@@
At inference time, given a target-user embedding and a reference sequence of eye movements, the trained DiffEyeSyn can inject the user-specific information of the target user into the reference eye movements (details in Section 4.5).
@@ 4.2. Conditional Diffusion Model for Eye Movement Synthesis
Furthermore, the inference process in Equation 4 is extended to:
$p_{\theta}\left(x_{t-1}\mid x_{t}\right)=\mathcal{N}\left(x_{t-1};\mu_{\theta}(x_{t},t\mid(x^{co}_{0},E(x_{0}))),\sigma_{\theta}(x_{t},t\mid(x^{co}_{0},E(x_{0})))I\right)$
@@ 4.3. User Identity Guidance
Training the model with $\mathcal{L}_{noise}$ allows the model to learn the distribution of human eye movements. However, there is no explicit user identity control in the generation process. Inspired by the classifier-guided diffusion model in image generation (Dhariwal and Nichol, 2021), we propose a novel user identity guidance that leverages the pretrained user authenticator $E$ to allow DiffEyeSyn to capture the user-specific information of human eye movements in the forward steps. At timestep $t$, we obtain $x_{t}$ by adding a sampled noise $\epsilon_{t}$ on the target $x_{0}$. We then predict the added noise $\hat{\epsilon_{t}}$ using Equation 6. With the predicted noise $\hat{\epsilon_{t}}$, we can directly estimate the target $\hat{x_{0}}$ by denoising $x_{t}$ using:
@@ 4.4. Data Preprocessing
1:Input: gaze data $q$, pretrained user authenticator $E$, total diffusion step $T$
@@
17:     $\nabla_{\theta}({\left\|{\epsilon_{t}-\hat{\epsilon_{t}}}\right\|}+\lambda(1-\frac{E(x_{0})\cdot E(\hat{x_{0}})}{|E(\hat{x_{0}})||E(\hat{x_{0}})|})$
@@ 4.5. Inference
A summary of DiffEyeSyn inference is shown in Algorithm 2. At the inference time, DiffEyeSyn injects the user-specific information into any given eye movements. It is worth noting that DiffEyeSyn also has the potential for user-specific gaze data imputation. This tasks aim to fill in missing values in existing eye movement sequences. To directly apply DiffEyeSyn at the inference time, users may utilise any interpolation technique to fill the missing values first, which, alongside a sequence of the target user’s eye movements, serves as input to DiffEyeSyn.
@@
1:Input: Any sequence preprocessed reference eye movement velocity $x^{co}$, pretrained user authenticator $E$, a sequence preprocessed eye movement velocity from the target user $x_{target-user}$, total number of denoising steps $T$
@@
5:Compute $\mu_{\theta}(x_{t},t\mid(x^{co},E(x_{\text{target-user}})))$
6:          and $\sigma_{\theta}(x_{t},t\mid(x^{co},E(x_{\text{target-user}})))$
7:          using $\hat{\epsilon_{t}}$
8:      Sample $x_{t-1}\sim p_{\theta}\left(x_{t-1}\mid x_{t}\right)$ using Equation 8
9:end for
10:return $x_{0}$
@@ 4.6. Architecture
$t_{\text{encoding }}=\left[\sin\left(10^{\frac{0\times 4}{63}}t\right),\cdots,\sin\left(10^{\frac{63\times 4}{63}}t\right),\cos\left(10^{\frac{0\times 4}{63}}t\right),\cdots,\cos\left(10^{\frac{63\times 4}{63}}t\right)\right]$
@@ 5.1. Experimental Setup
GazeBase is a large-scale dataset of eye movements collected using an Eyelink 1000 eye tracker that records at 1,000 Hz. It has 12,334 eye tracking records from 322 participants, covering different eye movements across nine rounds of recordings over 37 months. Each round has two sessions, and each session has seven visual tasks: fixations (FXS), horizontal (HSS) and random (RAN) saccades, reading (TEX), watching two videos (VD1 and VD2), and playing gaze-driven games (BLG). Further details on this dataset can be found in GazeBase paper (Griffith et al., 2021). To avoid the information leak, we adopted the dataset split described in the EKYT paper, as the EKYT model was trained on this dataset. In particular, the test set comprises recordings from 59 subjects presenting in round 6, while the training set comprises recordings from the remaining 263 participants. The recordings of the BLG task were excluded from the training set. Refer to the EKYT paper (Lohr and Komogortsev, 2022) for more details. There is no overlap of users between the training and test sets. These ensure zero information leakage from the pretrained model and we are able to test DiffEyeSyn performance on synthesising user-specific eye movements for unseen users during training.
@@
User Identity Removal. Previous studies in eye movement biometrics (Lohr and Komogortsev, 2022; Lohr et al., 2022) and gaze data super-resolution (Jiao et al., 2023) suggest that low-frequency gaze data at 20-30 Hz performs poorly in gaze-based user identification and authentication. In addition, Jiao et al. (Jiao et al., 2023) demonstrate that upsampling the low-frequency gaze data to a high frequency using traditional interpolation methods leads to a performance drop in user identification compared with the original low-frequency data. Based on these findings, we removed user-specific information in the ground truth gaze data by downsampling and interpolation. Specifically, we first downsampled the 1,000 Hz gaze data to 20 Hz and upsampled it back to 1,000 Hz using the previous interpolation. To verify the effectiveness of our identity removal, we calculated the cosine similarity between the EKYT embeddings extracted from the ground truth gaze data and the gaze data after the identity removal on both datasets. The resulting mean cosine similarity values on both datasets, as shown in Table 1, are consistently below $0.1$, confirming the effectiveness of our approach in removing user-specific information from the gaze data. Additionally, compared with adding random noise to the eye movements, removing the user-specific information by downsampling and interpolations keeps the low-frequency basis of the eye movements, which simplifies the task for DiffEyeSyn as it only needs to inject the user-specific subtle movements back to the given eye movement sequence, instead of learning to synthesis both realistic low-frequency and high-frequency eye movements at once.